# Sailor: Automating Distributed Training over Dynamic, Heterogeneous, and Geo-distributed Clusters

### v1

@@ 10 T_iter[P][i][R] = time_i
12 time_all = None
@@ 17 C_rem = Ci - cost_i
26 time_all = min(time_all, time_i)
27 if nextconf.straggler < assumed_straggler:
@@ 4.3. Sailor Simulator
The planner uses the simulator to evaluate the performance and memory footprint of the generated plans. The Sailor simulator takes as input a training job specification (model, global batch size, optimizer, hyperparameters) and a job parallelization plan. It then estimates the memory footprint per GPU, the iteration time, and the cost per iteration. The simulator allows the planner to easily specify different types of heterogeneity: hardware heterogeneity (GPU type, number of GPUs per node, and network bandwidth) and job configuration heterogeneity (number of pipelines and different stage configurations per pipeline). The training configuration also specifies the microbatch size. The simulator also incorporates the information collected by the profiler about the training job and network bandwidth of the used links.
@@
Memory footprint estimation: The Sailor simulator accurately estimates a training job’s memory footprint by: 1) calculating memory footprint per GPU, per-stage and 2) considering all main sources of memory footprint during training. Compared to prior works that assume a homogeneous memory footprint per stage, we observe that for a given parallelism configuration, the memory footprint of a training worker depends on layer partitioning, its pipeline stage index, its tensor parallelism degree, and its microbatch size. Hence, memory footprint varies among different workers and needs to be analyzed per worker to detect OOM scenarios.
@@
Second, the peak memory footprint of a worker throughout training consists of various sources, often ignored in prior works. The peak memory footprint $M_{peak}$ of a worker is given by: $M_{peak}=M_{model}+M_{activation}$, where $M_{model}$ corresponds to the memory needed to keep copies of model parameters and is given by $M_{model}=num\_params\cdot mul\_factor\cdot data\_type\_size$. $num\_params$ is found by the stage id and tensor parallelism degree of the worker, while $mul\_factor$ accounts for multiple copies needed for the model itself, the optimizer, gradients, and communication (Rajbhandari21Zero, ). $M_{activation}$ is the memory needed for storing layer activations and depends on the stage id and tensor parallelism degree of a worker, and microbatch size. By computing the memory footprint for each worker and comparing with the worker’s GPU capacity, the simulator can easily detect OOM cases.
@@
Iteration time estimation: We define one iteration as a full pass over the user-defined global batch size. The iteration time is calculated as: $T_{iter}=\max(T_{ppi})+T_{sync}+T_{update}$, where $T_{ppi}$ is the time needed for pipeline $i$, $T_{sync}$ is the time needed for gradient synchronization at the end of an iteration, and $T_{update}$ is the time for model update. We compute $T_{{pp}_{i}}$ and $T_{sync}$ following the formulas described in (strati24crossregion, ) for 1F1B pipeline parallelism, using our profiling for network bandwidth with respect to the message size per network link, to estimate communication time (for peer-to-peer and collectives). For each pipeline, the 1F1B pipeline parallelism schedule includes a warm-up, steady, and cool-down phase per iteration, where the steady phase is determined by the stage with the largest computation time (straggler). After computing the iteration time per pipeline, we compute synchronization time. Taking the maximum time per pipeline accounts for straggler effects caused by heterogeneity in GPU generations, inter-GPU, and CPU-GPU interconnects.
Iteration cost estimation: Related works (Table 1) do not compute the monetary cost of different resource allocation and job configuration combinations, since they only optimize for throughput. However, a very important metric to account for is cost per iteration, especially with geo-distributed training, due to costs associated with across-zone communication. Since Sailor does not change the global batch size and training hyperparameters, the number of iterations needed to reach convergence is constant, regardless of the cluster setup and parallelization strategy. Thus, the monetary cost per iteration indicates the total budget needed for the whole training. The metric depends both on the cost of allocated resources, as well as the training throughput and communication cost. The cost per iteration is given by $C_{iter}=C_{comp}+C_{comm}$, where $C_{comp}$ is the cost due to compute resources, and is calculated as $\sum_{i}({N_{i}\cdot cost\_per\_gpu_{i}})\cdot T_{iter}$, for all different GPU types $i$ in the cluster. $C_{comm}$ stands for the cost for data exchange per iteration (e.g., when using geo-distributed training in public cloud) and is given by $C_{comm}=\sum_{ij}(bytes_{ij}\cdot cost\_per\_byte\_{ij})$ for all zones $i,j$ in the cluster. $bytes_{ij}$ might include traffic for data and pipeline parallel communication.
@@ 4.4. Sailor Distributed Training Framework
The Sailor training framework receives the job configuration that the planner generates, sets up the cluster, and starts training. We modified Megatron-DeepSpeed (megatron-deepspeed, ) to support heterogeneous plans and seamless elasticity, to make it compatible to planner’s output and allow for fast reconfiguration when resource availability changes.
@@
Support for seamless fault-tolerance and elasticity: The Megatron-DeepSpeed framework lacks failure recovery and dynamic resource reconfiguration: the whole training needs to stop, and the user needs to manually reconfigure and restart the job. However, resource availability frequently changes both in the cloud (especially with spot instances) (Thorpe23Bamboo, ; strati24crossregion, ) and in on-premise datacenters as jobs start or finish (wagenlander24tenplex, ). We introduce modifications for fast, automated failure detection and reconfiguration, minimizing rollback time. Each job consists of a controller and multiple workers. The workers handle training, while the controller monitors their status and detects resource availability changes. Upon detecting a change, the controller reinvokes the planner to generate a new plan and instructs workers to adjust accordingly. We follow a kill-free approach to minimize the reconfiguration time, where the existing workers do not exit, but instead destroy the current communication group, clean up their GPU memory, repartition the model, and setup a new communication group. If additional resources become available, the controller waits for new workers to initialize before updating the training configuration. Training restarts from the latest available checkpoint. We use asynchronous checkpointing to minimize the rollback time (Mohan21CheckFreq, ; strati25pccheck, ).
@@ 5. Evaluation
System configurations: We evaluate Sailor using real hardware and simulations. We use different cluster setups and GPU generations, in both cloud environments and on-premise clusters. In the public cloud, we used VMs with A100-40GB and V100-16GB from GCP (gcpgpus, ). For our on-premise datacenter experiments, we used a cluster with up to 32 homogeneous machines with 4 Grace-Hopper GPUs each, and a cluster of heterogeneous machines with 2x8 Titan-RTX, 3x8 RTX-2080, and 2x8 RTX-3090.
Models: We use the OPT-350M (hfopt, ) and GPT-Neo-2.7B (hfgptneo, ) models, with global batch size of 2048 sequences, and sequence length of 2048 tokens, with the Adam optimizer.
Baselines: To our knowledge, we present the first comprehensive comparison of major planners for large-scale ML training, including planners targeting homogeneous resources (Piper (Tarnawski21Piper, ), AMP (li2022amp, ), Varuna (Athlur22Varuna, ), Aceso (Liu24Aceso, )), heterogeneous resources (Metis (Um24Metis, ), FlashFlex (yan2024flashflex, )), and geo-distributed training (DTFM(yuan2023decentralized, )). All baselines, except Aceso, are integrated into our platform with a unified Python API. We profile our models once and give each baseline its required profiling information. Aceso defines its own operators, which we profile separately. As Aceso also uses the Megatron backend, per-layer runtime profiles are very close to our models. We used the open-source version of all baselines. DTFM (yuan2023decentralized, ) does not determine parallelization strategies (e.g., DP, PP), but instead partitions a given plan. Therefore, we exhaustively generated all homogeneous parallelization plans and applied their partitioning methods to each. As the Atlas paper does not have an open-source implementation of runtime and memory simulation, we were unable to test Atlas end-to-end. However, we tested the zone assignment described in the paper (palak2024atlas, ), which performs similar to Sailor.
@@ 5.1. Validation of the Sailor simulator
Cluster of Heterogeneous GPU types. Figure 6 shows the iteration time prediction error for the OPT-350M model in a heterogeneous cluster of Titan-RTX, RTX-2080, and RTX-3090. The homogeneous planners (Piper, Varuna, Aceso) do not consider the differences in forward and backward passes of the different GPU types, resulting in an average error of 28%, 47%, and 37%, respectively. Even heterogeneous planners fail to accurately capture runtime: since FlashFlex relies on the theoretical performance of GPUs, it cannot accurately predict the runtime, getting an error of 69%. Metis fails to fully capture the heterogeneous network bandwidth between nodes, thus miscalculating the communication cost, resulting in 28% error in iteration time estimation, on average. In contrast, Sailor’s iteration time estimations error is 4.5% on average.
@@
We use our simulator to evaluate the throughput achieved by Sailor and the baselines across various cluster configurations. All baselines require a predefined resource topology as input: we consider 4-GPU VMs for each GPU type. Sailor takes resource quotas as input (total number of GPUs per type per zone) and jointly determines both the topology (VM allocation) and the parallelization plan. For Metis, we impose a 300-second time limit and use the best solution found within that period, if available. We summarize key takeaways.
@@ 5.2.1. Homogeneous Setups
Figure 7 shows the throughput achieved using the baseline planners and Sailor with only A100 GPUs for the OPT-350M model. Varuna failed to generate a valid plan that would not lead to OOM errors due to the poor memory estimation, and its limited search space (only supporting 2D parallelism). Sailor improves throughput by 1.15$\times$ compared to the closest baseline (DTFM), and even up to 5.7$\times$ (compared to Aceso).
5.2.2. Heterogeneous Setups
We evaluate Sailor’s throughput compared to baselines (AMP, Metis, and FlashFlex) in a heterogeneous cluster setup using a mixture of A100 and V100 for OPT-350M (Figure 8) and GPT-Neo-2.7B (Figure 9). We also compare the throughput achieved by Sailor, when using only homogeneous resources (either A100 or V100). We vary the ratio and amount of A100 and V100 (50%-50% and 25%-75%) to assess the impact of having more low-end GPUs in the cluster (guo2024cephalo, )). As in the homogeneous setup, all baselines get a fixed resource topology (4-GPU VMs) as input, while Sailor also determines the resource topology along with the parallelization plan. We also report the monetary cost per iteration at each scenario (number on top of bar), and the number of plans generated by the baseline that would lead to OOM before a valid plan was found (bold number on top of bar).
Throughput of heterogeneous baselines: Although AMP achieves high throughput in the homogeneous case, it performs poorly in the heterogeneous scenario, as it only allows for homogeneous plans, and does not correctly model stragglers. As a result, the monetary cost per iteration also increases. Furthermore, since AMP does not model training memory footprint, it leads to a large number of OOM plans, especially in the case of the GPT-Neo model (Figure  9).
Sailor achieves significantly higher throughput compared to baselines: for the OPT-350M model, Sailor achieves 1.9$\times$, 2.03$\times$, and 1.15$\times$ higher throughput compared to AMP, FlashFlex, and Metis, respectively, when the ratio of A100 and V100 GPUs is equal, and 1.57$\times$, 1.55$\times$, 1.39$\times$ higher throughput when more V100 are available. The speedups are similar for the GPT-Neo model as well. Also, Sailor does not output invalid plans, significantly improving the plan deployment compared to baselines with 10s of invalid plans. Sailor’s ability to discover efficient resource topologies and parallelization plans also translates to significantly lower cost compared to the baselines (up to 2.67$\times$ lower cost compared to baselines in Figures 8 and 9).
Search times: Table 2 shows the search time for Figure 9(b) for the different baselines. Metis is capped by 300 sec in all cases, after when its search is interrupted. The other baselines are significantly faster, as they manage to finish the search in less than 200 sec for all cases. Sailor keeps the search time within 1 minute even in the largest cluster case of 512 GPUs, which demonstrates its efficient searching algorithm.
@@ 5.2.2. Heterogeneous Setups
Benefits of heterogeneity: Figures 8 and 9 also show the throughput achieved by Sailor when given homogeneous-only setups of A100 (Sailor-A100) or V100 GPUs (Sailor-V100). Given that V100s are less efficient than A100s, using A100 only, or a mixture of A100 and V100 is always better than using V100 only. However, using V100 in addition to A100 does not always improve throughput. Heterogeneity is more beneficial when resources are limited (e.g. Figure 8(a), with 32 GPUs per type), or with larger models like GPT-Neo. In fact, for the OPT-350M model, when 128 A100 and 128 V100 are available, Sailor chooses a plan with 128 A100 only, as it determines that no additional benefits will be gained by adding extra resources. Moreover, when the ratio of V100 to A100 is higher than 1, the throughput improvents are more significant, as shown in Figures 8(b) and 9(b), where the V100:A100 ratio is 3:1. This aligns with the GPU memory capacity ratio, as well as the time for forward/backward pass for the transformer layers of the two GPU types, enabling better load balancing. Finally, heterogeneity leads to a higher cost, as more resources are used. Since the objective is to maximize throughput, Sailor ignored monetary cost when searching job configurations. In 5.2.4, we will show Sailor’s ability to consider budget optimization and constraints.
@@
Table 2. Search times (in sec) for Figure 9(b).

### v2

@@ 10 T_iter[P][i][R] = time_i
12 time_all = None, min_time = None
@@ 17 C_rem = Ci - cost_i
27 min_time = min(time_all, min_time)
@@ 4.3. Sailor Simulator
The planner uses the simulator to evaluate the performance and memory footprint of the generated plans. The Sailor simulator takes as input a training job specification (model, global batch size, optimizer, hyperparameters) and a job parallelization plan. It then estimates the memory footprint per GPU, the iteration time, and the cost per iteration. The simulator allows the planner to easily specify different types of heterogeneity: hardware heterogeneity (GPU type, number of GPUs per node, and network bandwidth) and job configuration heterogeneity (number of pipelines and different stage configurations per pipeline). The training configuration also specifies the microbatch size. The simulator also incorporates the information collected by the profiler about the training job and network bandwidth of the used links. Both the Sailor planner and simulator treat GPUs as black-box compute units, thus they can seamlessly support GPUs from different generations, vendors, and even different accelerators (e.g. TPUs).
@@
Memory footprint estimation: The Sailor simulator accurately estimates a training job’s memory footprint by: 1) calculating memory footprint per GPU, per-stage and 2) considering all main sources of memory footprint during training. Compared to prior works that assume a homogeneous memory footprint per stage, we observe that for a given parallelism configuration, the memory footprint of a training worker depends on its layer partitioning, pipeline stage index, tensor parallelism degree, and microbatch size. Hence, memory footprint varies among workers and needs to be analyzed per worker to detect OOM scenarios.
@@
Second, the peak memory footprint of a worker throughout training consists of various sources, often ignored in prior works. The peak memory footprint $M_{peak}$ of a worker is given by: $M_{peak}=M_{model}+M_{activation}$, where $M_{model}$ corresponds to the memory needed to keep copies of model parameters and is given by $M_{model}=num\_params\cdot mul\_factor\cdot data\_type\_size$. num_paramsnum\_params is found by the stage id and tensor parallelism degree of the worker, while mul_factormul\_factor accounts for multiple copies needed for the model itself, the optimizer, gradients, and communication (Rajbhandari et al., 2021). $M_{activation}$ is the memory needed for storing layer activations and depends on the stage id and tensor parallelism degree of a worker, and microbatch size. By computing the memory footprint for each worker and comparing with the worker’s GPU capacity, the simulator can easily detect OOM cases.
@@
Iteration time estimation: We define one iteration as a full pass over the user-defined global batch size. The iteration time is calculated as: $T_{iter}=\max(T_{ppi})+T_{sync}+T_{update}$, where $T_{ppi}$ is the time needed for pipeline $i$, $T_{sync}$ is the time needed for gradient synchronization at the end of an iteration, and $T_{update}$ is the time for model update. We compute $T_{{pp}_{i}}$ and $T_{sync}$ following the formulas described in (Strati et al., 2024a) for 1F1B pipeline parallelism, using our profiling for network bandwidth with respect to the message size per network link, to estimate communication time (for peer-to-peer and collectives). For each pipeline, the 1F1B pipeline parallelism schedule includes a warm-up, steady, and cool-down phase per iteration, where the steady phase is determined by the stage with the largest computation time (straggler). After computing the iteration time per pipeline, we compute synchronization time. Taking the maximum time per pipeline accounts for straggler effects caused by heterogeneity in GPU generations, inter-GPU, and CPU-GPU interconnects.
[…]
@@ 4.4. Sailor Distributed Training Framework
The Sailor training framework receives the job configuration that the planner generates, sets up the cluster, and starts training. We modified Megatron-DeepSpeed (deepspeedai, 2025) to support heterogeneous plans and seamless elasticity, to make it compatible to planner’s output and allow for fast reconfiguration when resource availability changes.
@@
Support for fault-tolerance and elasticity: The Megatron-DeepSpeed framework lacks failure recovery and dynamic resource reconfiguration: the whole training needs to stop, and the user needs to manually reconfigure and restart the job. However, resource availability frequently changes both in the cloud (especially with spot instances) (Thorpe et al., 2023; Strati et al., 2024a) and in on-premise datacenters as jobs start or finish (Wagenländer et al., 2024). We introduce modifications for fast reconfiguration. Each job consists of a controller and multiple workers. The workers handle training, while the controller monitors their status and detects resource availability changes. Upon detecting a change, the controller reinvokes the planner to generate a new plan and instructs workers to adjust accordingly. We follow a kill-free approach to minimize reconfiguration time: existing workers destroy the current communication group, clean up their GPU memory, repartition the model, and setup a new communication group. If additional resources become available, the controller waits for new workers to initialize before updating the training configuration. Training restarts from the latest available checkpoint. We use asynchronous checkpointing to minimize the rollback time (Mohan et al., 2021; Strati et al., 2025).
@@ 5. Evaluation
System configurations: We evaluate Sailor using real hardware and simulations. We use different cluster setups and GPU generations, in both cloud environments and on-premise clusters. In the public cloud, we used VMs with A100-40GB and V100-16GB from GCP (Cloud, 2024). For our on-premise datacenter experiments, we used a cluster with up to 32 homogeneous machines with 4 Grace-Hopper GPUs each, and a cluster of heterogeneous machines with 2x8 Titan-RTX, 3x8 RTX-2080, and 2x8 RTX-3090.
Models: We use the OPT-350M (HuggingFace, 2025b) and GPT-Neo-2.7B (HuggingFace, 2025a) models, with global batch size of 2048 sequences, and sequence length of 2048 tokens, with the Adam optimizer.
Baselines: To our knowledge, we present the first comprehensive comparison of major planners for large-scale ML training, including planners targeting homogeneous resources (Piper (Tarnawski et al., 2021), Varuna (Athlur et al., 2022), Aceso (Liu et al., 2024)), heterogeneous resources (AMP (Li et al., 2022), Metis (Um et al., 2024), FlashFlex (Yan et al., 2024)), and geo-distributed training (DTFM(Yuan et al., 2023)). All baselines, except Aceso, are integrated into our platform with a unified Python API. We profile our models once and give each baseline its required profiling information. Aceso defines its own operators, which we profile separately. As Aceso also uses the Megatron backend, per-layer runtime profiles are very close to our models. We used the open-source version of all baselines. DTFM (Yuan et al., 2023) does not determine parallelization strategies (e.g., DP, PP), but instead partitions a given plan. Therefore, we exhaustively generated all homogeneous parallelization plans and applied their partitioning methods to each. As the Atlas paper does not have an open-source implementation of runtime and memory simulation, we were unable to test Atlas end-to-end. However, we tested the zone assignment described in the paper (Palak et al., 2024a), which performs similar to Sailor.
@@ 5.1. Validation of the Sailor simulator
Cluster of Heterogeneous GPU types. Figure 6 shows the iteration time prediction error for the OPT-350M model in a heterogeneous cluster of Titan-RTX, RTX-2080, and RTX-3090. The homogeneous planners (Piper, Varuna, Aceso) do not consider the differences in forward and backward passes of the different GPU types, resulting in an average error of 28%, 47%, and 37%, respectively. Even heterogeneous planners fail to accurately capture runtime: since FlashFlex relies on the theoretical performance of GPUs, it cannot accurately predict the runtime, getting an error of 69%. Metis fails to fully capture the heterogeneous network bandwidth between nodes, thus miscalculating the communication cost, resulting in 28% error in iteration time estimation, on average.On average, Sailor’s iteration time estimations error is 4.5%.
@@
We evaluate the throughput achieved by Sailor and the baselines across various cluster configurations using both real hardware and our simulator. All baselines require a predefined resource topology as input: we consider 4-GPU VMs for each GPU type. Sailor takes resource quotas as input (total number of GPUs per type per zone) and jointly determines both the topology (VM allocation) and the parallelization plan. For Metis, we impose a 300-second time limit and use the best solution found within that period, if available. We summarize key takeaways.
@@ 5.2.1. Homogeneous Setups
Figure 7 shows the throughput achieved using the baseline planners and Sailor with only A100 GPUs for the OPT-350M model. Varuna failed to generate a valid plan that would not lead to OOM errors due to the poor memory estimation, and its limited search space (only supporting 2D parallelism). Sailor improves throughput by 1.15 compared to the closest baseline (DTFM), and even up to 5.7 (compared to Aceso).
@@
5.2.2. Heterogeneous Setups
We evaluate Sailor’s throughput compared to baselines (AMP, Metis, and FlashFlex) in a heterogeneous cluster setup using a mixture of A100 and V100. We vary the ratio and amount of A100 and V100 (50%-50% and 25%-75%) to assess the impact of having more low-end GPUs in the cluster (Guo et al., 2024)). As in the homogeneous setup, all baselines get a fixed resource topology (4-GPU VMs) as input, while Sailor also determines the resource topology along with the parallelization plan.
Throughput of heterogeneous baselines: Figures 8 and 9 use our simulator to evaluate Sailor and the baselines in large heterogeneous clusters for the OPT-350M and GPT-Neo-2.7B model, respectively. We also compare the throughput achieved by Sailor, when using only homogeneous resources (either A100 or V100). We also report the monetary cost per iteration at each scenario (number on top of bar), and the number of plans generated by the baseline that would lead to OOM before a valid plan was found (bold number on top of bar). Although AMP achieves high throughput in the homogeneous case, it performs poorly in the heterogeneous scenario, as it only allows for homogeneous plans, and does not correctly model stragglers. As a result, the monetary cost per iteration also increases. Furthermore, since AMP does not model training memory footprint, it leads to a large number of OOM plans, especially in the case of the GPT-Neo model (Figure  9).
Sailor achieves significantly higher throughput compared to baselines: for the OPT-350M model, Sailor achieves 1.9, 2.03, and 1.15 higher throughput compared to AMP, FlashFlex, and Metis, respectively, when the ratio of A100 and V100 GPUs is equal, and 1.57, 1.55, 1.39 higher throughput when more V100 are available. The speedups are similar for the GPT-Neo model as well. Compared to baselines, Sailor uses larger tensor parallelism and longer pipelines, accounting for data parallelism limitations among heterogeneous nodes (for example AMP uses data parallelism of 256 in Figure 12 which significantly increases the data parallelism communication cost.) Also, Sailor does not output invalid plans, significantly improving the plan deployment compared to baselines with 10s of invalid plans. Sailor’s ability to discover efficient resource topologies and parallelization plans also translates to significantly lower cost compared to the baselines (up to 2.67 lower cost compared to baselines in Figures 8 and 9).
We also evaluated Sailor and the other heterogeneous baselines with real hardware using a smaller cluster of A100 and V100 GPUs. Figure 10 shows the throughput achieved for the OPT-350M model when using an equal amount of GPUs per type (8 each), and when using more V100 than A100, as V100 were more readily available. When using the same number of GPUs per type, Sailor outperforms the baselines by 1.08-1.81 and does not generate any Out-Of-Memory plans. In contrast, AMP and Metis generate 5 and 1 invalid (OOM) plans before finding a valid plan, respectively. The valid plan found by Metis is similar to Sailor’s, but Sailor avoids invalid outputs entirely, and finds a solution in less than 1 sec, while Metis takes 60 sec. When using 8 A100 and 16 V100 GPUs, Metis fails to output a plan as it requires the global batch size to be equally divisible by the total number of GPUs. We therefore reuse the plan from the 16 GPU case. AMP produces the same plan for the 24 GPU case as the 16 GPU case, while the plan provided by FlashFlex utilizes all 24 GPUs but uses an unnecessarily large number of pipeline stages that degrades throughput. Sailor outperforms the baselines by 1.19-2 for the 24 GPU scenario. Our simulator’s estimated iteration times were within 4% of those measured on real hardware.
Search times: Table 2 shows the search time for Figure 9(b) for the different baselines. Metis has long search times. In our experiments, we let Metis search for up to 300 seconds and take the best plan it produces in this time window. The other baselines are significantly faster, finishing their search in less than 200 seconds in all cases. Sailor keeps the search time within 1 minute even for the largest case (512 GPUs) due to its efficient search algorithm.
@@ 5.2.2. Heterogeneous Setups
Benefits of heterogeneity: Figures 8, 9 and 10 also show the throughput achieved by Sailor when given homogeneous-only setups of A100 (Sailor-A100) or V100 GPUs (Sailor-V100). Given that V100s are less efficient than A100s, using A100 only, or a mixture of A100 and V100 is always better than using V100 only. However, using V100 in addition to A100 does not always improve throughput. Heterogeneity is more beneficial when resources are limited (e.g. Figure 8(a), with 32 GPUs per type), or with larger models like GPT-Neo. In fact, for the OPT-350M model, when 128 A100 and 128 V100 are available, Sailor chooses a plan with 128 A100 only, as it determines that no additional benefits will be gained by adding extra resources. Moreover, when the ratio of V100 to A100 is higher than 1, the throughput improvents are more significant, as shown in Figures 10, 8(b) and 9(b), where the V100:A100 ratio is 2:1 or 3:1. This aligns with the GPU memory capacity ratio, as well as the time for forward/backward pass for the transformer layers of the two GPU types, enabling better load balancing. Finally, heterogeneity leads to a higher cost, as more resources are used. Since the objective is to maximize throughput, Sailor ignored monetary cost when searching job configurations. In §5.2.4, we will show Sailor’s ability to consider budget optimization and constraints.
@@
Table 2. Search times (in seconds) for Figure 9(b).
Figure 10. Throughput of heterogeneous planners for clusters of A100 and V100 GPUs for the OPT-350M model.